Hultquist Corporation has two manufacturing departments--Forming and Customizing. The company used the following data at the beginning of the period to calculate predetermined overhead rates:

Forming Customizing Total
Estimated total machine-hours (MHs) 2,000 8,000 10,000
Estimated total fixed manufacturing overhead cost $ 6,400 $ 20,800 $ 27,200
Estimated variable manufacturing overhead cost per MH $ 1.50 $ 3.00

During the period, the company started and completed two jobs--Job C and Job L. Data concerning those two jobs follow:

Job C Job L
Direct materials $ 13,900 $ 7,800
Direct labor cost $ 21,000 $ 7,800
Forming machine-hours 1,250 750
Customizing machine-hours 1,250 6,750

Required:

a. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate that overhead rate. (Round your answer to 2 decimal places.)

b. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate the amount of manufacturing overhead applied to Job L. (Do not round intermediate calculations.)

c. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate the total manufacturing cost assigned to Job L. (Do not round intermediate calculations.)

d. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours and uses a markup of 20% on manufacturing cost to establish selling prices. Calculate the selling price for Job L. (Do not round intermediate calculations.)

e. Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments. What is the departmental predetermined overhead rate in the Forming department? (Round your answer to 2 decimal places.)

f. Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both production departments. What is the departmental predetermined overhead rate in the Customizing department? (Round your answer to 2 decimal places.)

g. Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both production departments. How much manufacturing overhead will be applied to Job L? (Do not round intermediate calculations.)

h. Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both production departments. Further assume that the company uses a markup of 20% on manufacturing cost to establish selling prices. Calculate the selling price for Job L. (Do not round intermediate calculations.)
Answer #1

a) Overhead rate = (2000*1.5+8000*3+27200)/10000 = 5.42 per hour

b) Overhead applied = (750+6750)*5.42 = $40650

c) Total manufacturing cost = 7800+7800+40650 = 56250

d) Bid price = 56250*1.2 = 67500

e) Forming overhead rate = (6400/2000)+1.50 = 4.70 per hour

f) Customizing overhead rate = (20800/8000)+3 = 5.60 per hour

g) Overhead applied = (750*4.7+6750*5.60) = 41325

h) Total manufacturing cost = 7800+7800+41325 = 56925

Selling price = 56925*1.2 = $68310